OpenStudy (anonymous):

Solve each equation. 5x2 + 500 = 844.45

OpenStudy (anonymous):

is the x serving as a multiplication sign or a variable? @amberkels55

OpenStudy (anonymous):

Variable!

OpenStudy (anonymous):

5x^2

OpenStudy (anonymous):

Well first you need to isolate the variable so how would you begin doing that

OpenStudy (anonymous):

@amberkels55

OpenStudy (yanasidlinskiy):

Solving 500 + 5x^2 = 844.45 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-500' to each side of the equation. 500 + -500 + 5x^2 = 844.45 + -500 Combine like terms: 500 + -500 = 0 0 + 5x^2 = 844.45 + -500 5x^2 = 844.45 + -500 Combine like terms: 844.45 + -500 = 344.45 5x^2 = 344.45 Divide each side by '5'. x^2 = 68.89 Simplifying x^2 = 68.89 Take the square root of each side: x = {-8.3, 8.3}

OpenStudy (anonymous):

444444545454545454545445454545454545454545454545454545454545454545454545454545454545454545454454545453554455445544555444554445555444555454545454545454545454545454545454545454545454